Please work it out so that I can understand. Thanks.

Island Novelties, Inc., of Palau makes two products—Hawaiian Fantasy and Tahitian Joy. Each product’s selling price, variable expense per unit, and annual sales volume are as follows:

Hawaiian Fantasy Tahitian Joy
Selling price per unit $ 20 $ 100
Variable expense per unit $ 13 $ 30
Number of units sold annually 34,000 7,200

Fixed expenses total $651,900 per year.

Required:

1. Assuming the sales mix given above, do the following:

a. Prepare a contribution format income statement showing both dollar and percent columns for each product and for the company as a whole.

b. Compute the company's break-even point in dollar sales. Also, compute its margin of safety in dollars and its margin of safety percentage.

2. The company has developed a new product called Samoan Delight that sells for $50 each and that has variable expenses of $30 per unit. If the company can sell 12,000 units of Samoan Delight without incurring any additional fixed expenses:

a. Prepare a revised contribution format income statement that includes Samoan Delight. Assume that sales of the other two products does not change.

b. Compute the company’s revised break-even point in dollar sales. Also, compute its revised margin of safety in dollars and margin of safety percentage.
Answer #1
Dear Student
Thank you for using homeworklib
Please find below the answer   
Statementshowing Computations Hawaiian Fantasy Tahitian Joy Total
Paticulars Amount % Amount % Amount %
Sales            680,000.00 100.00%               720,000.00 100.00%       1,400,000.00 100.00%
Less Variable Expenses            442,000.00 65.00%               216,000.00 30.00%          658,000.00 47.00%
Contribution Margin            238,000.00 35.00%               504,000.00 70.00%          742,000.00 53.00%
Fixed cost 0.00% 0.00%          651,900.00 46.56%
Net operating income 0.00% 0.00%             90,100.00 6.44%
Contribution Margin Ratio 53.00%
BEP in Sales $ = 651,900/53%       1,230,000.00
MOS in $ = 1400,000 - 1230,000            170,000.00
MOS in % = (170,000)/1400,000 12.14%
2)
Statementshowing Computations Hawaiian Fantasy Tahitian Joy Samoan Delight Total
Paticulars Amount % Amount % Amount % Amount %
Sales            680,000.00 100.00%               720,000.00 100.00%          600,000.00 100.00%    2,000,000.00 100.00%
Less Variable Expenses            442,000.00 65.00%               216,000.00 30.00%          360,000.00 60.00%    1,018,000.00 50.90%
Contribution Margin            238,000.00 35.00%               504,000.00 70.00%          240,000.00 40.00%        982,000.00 49.10%
Fixed cost        651,900.00 32.60%
Net operating income        330,100.00 16.51%
Contribution Margin Ratio 49.10%
BEP in Sales $ = 651,900/53%    1,327,698.57
MOS in $ = 2000,000-1327,698.57            672,301.43
MOS in % = 672,301.43/2000,000 33.62%